22west:

How many triangles can be constructed with angles measuring 60º, 60º, and 40º? A. one B. none C. more than one

22west:

@dude

dude:

Same as previous question, 1 triangle must be 180 degrees

vaporeon246:

60+ 60 =120 + 40 = 160

22west:

So it won't equal even one.

dude:

Yup